In 2021, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2023. Information related to the contract is as follows:

2021 2022 2023
Cost incurred during the year $ 2,640,000 $ 2,300,000 $ 2,926,000
Estimated costs to complete as of year-end 6,160,000 2,660,000 0
Billings during the year 2,080,000 2,860,000 5,060,000
Cash collections during the year 1,840,000 2,800,000 5,360,000


Westgate recognizes revenue over time according to percentage of completion.

Problem 6-10 (Algo) Part 5

5. Calculate the amount of revenue and gross profit (loss) to be recognized in each of the three years assuming the following costs incurred and costs to complete information. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount. Loss amounts should be indicated with a minus sign.)

2021 2022 2023
Costs incurred during the year $ 2,640,000 $ 3,840,000 $ 4,020,000
Estimated costs to complete as of year-end 6,160,000 4,180,000 0
Answer #1

Answers

  • Requirement 5

2021

2022

2023

Revenue

$                3,000,000.00

$               3,078,799

$                    3,921,201

Gross Profit (Loss)

$                   360,000.00

$             (1,020,000.00)

$                        160,000.00

  • Workings

Working

2021

2022

2023

A

Contract Price

$             10,000,000.00

$             10,000,000.00

$                  10,000,000.00

B

Cost Incurred to Date

$                2,640,000.00

$               6,480,000.00

$                  10,500,000.00

C

Estimated cost yet to be incurred to complete the contract

$                6,160,000.00

$               4,180,000.00

$                                         -  

D = B+C

Total Cost

$                8,800,000.00

$             10,660,000.00

$                  10,500,000.00

E = (B/D) x 100

% of Completion

30.00%

60.79%

100.00%

F = A x E

Revenue to date

$                3,000,000.00

$               6,078,799.25

$                  10,000,000.00

G

Revenue of Previous year

$                                     -  

$               3,000,000.00

$                    6,078,799.25

H = F - G

Net Revenue this year

$                3,000,000.00

$               3,078,799.25

$                    3,921,200.75

I (=B)

Cost to date

$                2,640,000.00

$               6,480,000.00

$                  10,500,000.00

J

Cost to date of previous year

$                                     -  

$               2,640,000.00

$                   6,480,000.00

K = I - J

Net Cost for the year

$                2,640,000.00

$               3,840,000.00

$                    4,020,000.00

L = A - D

Estimated Gross Profit (Loss)

$                1,200,000.00

$                 (660,000.00)

$                     (500,000.00)

M

Gross Profit (Loss) to be recognised

$                   360,000.00

$             (1,020,000.00)

$                        160,000.00